A bag of peanuts is worth $0.24 less than a bag of cashews. Equal amounts of peanuts and cashews are used to make 44 bags of a mixture that sells for $1.45 per bag. How much is a bag of cashews worth?

Answer by ankor@dixie-net.com(22740)

"A bag of peanuts is worth $0.24 less than a bag of cashews."
Let x = cost of cashews
Then
(x-.24) = cost of peanuts
:
"Equal amounts of peanuts and cashews are used to make 44 bags of a mixture that sells for $1.45 per bag."
:
Since they say equal amts of both nuts, lets write a bag/cost equation
:
22b cashw + 22b peants = 44b mixture
22x + 22(x-.24) = 1.45(44)
:
22x + 22x - 5.28 = 63.80
:
22x + 22x = 63.80 + 5.28
:
44x = 69.08
:
x = 69.08/44
:
x = $1.57 cost per bag are the cashews
:
:
Check solution:
Peanuts cost: 1.57 - .24 = 1.33
1.57(22) + 1.33(22) = 1.45(44)
34.54 + 29.26 = 63.80; confirms our solution